# A Comparative Study of Online Disinformation and Offline Protests

### v4

@@ Abstract
In early 2021 the United States Capitol in Washington was stormed during a riot and violent attack. A similar storming occurred in Brazil in 2023. Although both attacks were instances in longer sequences of events, these have provided a testimony for many observers who had claimed that online actions, including the propagation of disinformation, have offline consequences. Soon after, a number of papers have been published about the relation between online disinformation and offline violence, among other related relations. Hitherto, the effects upon political protests have been unexplored. This paper thus evaluates such effects with a time series cross-sectional sample of $125$ countries in a period between 2000 and 2019. The results are mixed. Based on Bayesian multi-level regression modeling, (i) there indeed is an effect between online disinformation and offline protests, but the effect is partially meditated by political polarization. The results are clearer in a sample of countries belonging to the European Economic Area. With this sample, (ii) offline protest counts increase from online disinformation disseminated by domestic governments, political parties, and politicians as well as by foreign governments. Furthermore, (iii) Internet shutdowns tend to decrease the counts, although, paradoxically, the absence of governmental online monitoring of social media tends to also decrease these. With these results, the paper contributes to the blossoming disinformation research by modeling the impact of disinformation upon offline phenomenon. The contribution is important due to the various policy measures planned or already enacted.
@@ 1 Introduction
Recently, Piazza, (2021) published an intriguing study on the relationship between online disinformation and domestic terrorism. According to his result, disinformation spread by governmental actors and political parties contributes to domestic terrorism but the effect is mediated by political polarization and tribalism within countries. Rather similar results were recently obtained also by Gallacher et al., (2021) according to whom discussions in social media between opposing groups are associated with offline physical violence. Also Arayankalam and Krishna, (2021) showed that disinformation spread by foreign actors is associated with physical violence. A relation has been also found between online hate speech and offline hate crimes (Williams et al., 2020). Recently, Diab et al., (2023) found that white supremacist online propaganda tends to later appear offline in the form of fliers, banners, and graffiti. Closer to the present study, Kirkizh and Koltsova, (2021) further found that exposure to online news is positively associated with participation in political demonstrations, including particularly in authoritarian countries. While there is thus a solid foundation for the present work, thus far, the relation between online disinformation and offline political protests has not been explored. This gap in existing research is surprising already due to the studies exploring the relation of online phenomena to violence, which is a much stronger presumption than a mere political protest.
@@ 2 Framing
The disinformation and propaganda disseminated by ISIS and Russia have had one thing common; the primary (but not the only) target has been Western democracies, and the intention has been to deepen the existing societal cleavages and sabotage the internal cohesion within Western countries by exploiting weaknesses in the current engagement economy and conducting violent offline attacks, whether via terrorism or assassinations (Dundon and Houck, 2022; Hamilton, 2019; Kalniete and Pildegovic̆s, 2021; Rosenblatt et al., 2019; Sultan, 2019; Zhang et al., 2021). Hate is a powerful method, whether in the hands of Muslim minorities or right-wing extremists.
To reach the goals, which, as said, are a distinct trait of disinformation, time and sufficient resources are needed. In a sense, one still metaphorically needs the printing press, the airplane, or the television channel; to exploit societal weaknesses, one needs to be well-educated. The same applies when the goal of the exploitation is to provoke political protests in foreign countries. Education, staffing, and resources are thus another factor separating state-centric disinformation from other false information propagated by laymen. Though: if the online and offline realms intervene, so do the old logic and the new logic of media; many national disinformation outlets, some of which are linked to foreign disinformation organizations, rely on national media for their criticism and distortion of narratives (Toivanen et al., 2021). For building such disinformation outlets, nevertheless, time is needed even when the result would be amateurish due to lack of education.
@@ 3.3 Independent Variables
Six independent variables are included for capturing the policy-side. A few examples are in order to contextualize this side. Although heavily criticized (Schulz, 2020), the Network Enforcement Act (NetzDG) in Germany is a good example of legislative attempts to filter online hate speech. Likewise, in 2018, a comparable law was passed in France; it allows public authorities to remove fake information or even block sites that publish such content (Nagasako, 2020). In addition to these national laws, the European Union has been actively enacting different legislative measures and informal codes of conduct, including for disinformation and hate speech (Buiten et al., 2020; Pataki et al., 2021). On the side of even more drastic measures, particularly authoritarian countries in conflict zones have increasingly resorted to Internet shutdowns; among the memorable historical examples was the 2011 shutdown of the Internet in Egypt in the face of the Arab Spring uprisings (Arnaudo et al., 2013). Thus, there is a well-founded rationale to consider also the policy-side variables.
@@ 4 Results
Intercept $-2.59$ $-4.69$ $-0.43$ $-2.69$ $-4.83$ $-0.53$ $-2.91$ $-4.92$ $-0.88$
Gov. disinformation: domestic $-0.17$ $-0.26$ $-0.09$ $-0.10$ $-0.19$ $0.00$
Gov. disinformation: abroad $-0.11$ $-0.21$ $-0.01$ $-0.10$ $-0.21$ $0.02$
Foreign gov. disinformation $-0.07$ $-0.16$ $0.02$ $-0.07$ $-0.16$ $0.02$
Party disinformation: domestic $-0.13$ $-0.21$ $-0.05$ $0.02$ $-0.07$ $0.10$
Party disinformation: abroad $-0.10$ $-0.18$ $0.00$ $0.06$ $-0.05$ $0.16$
@@
Gov. filtering practice $0.05$ $-0.05$ $0.15$
Gov. shutdown capacity $-0.18$ $-0.28$ $-0.07$
Gov. social media alternatives $0.03$ $-0.10$ $0.16$
Gov. social media monitoring $-0.20$ $-0.29$ $-0.10$
Regulation of online content $0.09$ $-0.01$ $0.19$
@@
National online media $0.01$ $-0.05$ $0.07$ $0.01$ $-0.05$ $0.07$ $-0.02$ $-0.08$ $0.05$
Regime change $-0.03$ $-0.13$ $0.06$ $-0.01$ $-0.11$ $0.08$ $0.01$ $-0.09$ $0.11$
@@
Intercept $-4.87$ $-14.05$ $3.25$ $-4.33$ $-12.22$ $2.87$ $-2.55$ $-14.12$ $7.75$
Gov. disinformation: domestic $0.13$ $-0.06$ $0.32$ $0.47$ $0.17$ $0.76$
Gov. disinformation: abroad $-0.74$ $-1.04$ $-0.43$ $-0.61$ $-0.99$ $-0.21$
@@
Party disinformation: domestic $0.07$ $-0.11$ $0.25$ $0.61$ $0.34$ $0.88$
Party disinformation: abroad $-0.80$ $-1.06$ $-0.56$ $-0.62$ $-0.96$ $-0.28$
Gov. filtering capacity $0.12$ $-0.34$ $0.59$
Gov. filtering practice $-0.51$ $-0.94$ $-0.06$
Gov. shutdown capacity $-0.05$ $-0.56$ $0.48$
@@
Gov. social media monitoring $-0.38$ $-0.65$ $-0.10$
Regulation of online content $-0.38$ $-0.65$ $-0.10$
@@
Regime change $-0.28$ $-3.65$ $3.10$ $-1.08$ $-3.94$ $1.79$ $-2.15$ $-6.52$ $2.15$
$\ln(\textmd{population})$ $0.47$ $0.00$ $1.00$ $0.59$ $0.19$ $1.05$ $0.45$ $-0.16$ $1.14$
$\ln(\textmd{GDP per capita})$ $-0.15$ $-0.45$ $0.15$ $-0.36$ $-0.66$ $-0.07$ $-0.28$ $-0.60$ $0.03$
@@
Intercept $-0.98$ $-4.84$ $1.79$ $-18.46$
Gov. disinformation: domestic $0.41$ $-0.35$ $-0.14$ $0.24$
Gov. disinformation: abroad $-0.47$ $0.04$ $0.58$ $-0.48$
Foreign gov. disinformation $0.53$ $-0.28$ $<0.01$ $0.53$
Party disinformation: domestic $-0.47$ $0.16$ $-1.25$ $0.48$
Party disinformation: abroad $0.51$ $-0.29$ $-1.06$ $-0.98$
Gov. filtering capacity $-0.06$ $0.15$ $0.15$ $-0.69$
Gov. filtering practice $-0.18$ $0.17$ $0.68$ $0.07$
Gov. shutdown capacity $-0.95$ $-0.03$ $0.32$ $0.82$
Gov. social media alternatives $0.29$ $-0.04$ $-0.65$ $0.55$
Gov. social media monitoring $-0.60$ $-0.07$ $-0.21$ $-0.38$
Regulation of online content $-0.71$ $0.36$ $-0.25$ $-0.45$
National online media $0.16$ $-0.03$ $-0.11$ $-0.01$

### v5

@@ Abstract
This paper evaluates the effect of online disinformation upon offline political protests with a time series cross-sectional sample of $125$ countries in a period between 2000 and 2019. The results are mixed. Based on Bayesian multi-level regression modeling, (i) there indeed is an effect between online disinformation and offline protests, but the effect is partially meditated by political polarization. The results are clearer in a sample of countries belonging to the European Economic Area. With this sample, (ii) offline protest counts increase from online disinformation disseminated by domestic governments, political parties, and politicians as well as by foreign governments. Furthermore, (iii) Internet shutdowns tend to decrease the counts, although, paradoxically, the absence of governmental online monitoring of social media tends to also decrease these. With these results, the paper contributes to the blossoming disinformation research by modeling the impact of disinformation upon offline phenomenon. The contribution is important due to the various policy measures planned or already enacted.
@@ 1 Introduction
In early 2021 the United States Capitol in Washington was stormed during a riot and violent attack. A similar storming occurred in Brazil in 2023. Although both attacks were instances in longer sequences of events, these have provided a testimony for many observers who had claimed that online actions, including the propagation of disinformation, have offline consequences. Soon after, a number of papers have been published about the relation between online disinformation and offline violence, among other related relations.
Recently, an intriguing study was published on the relationship between online disinformation and domestic terrorism. According to the paper’s result, disinformation spread by governmental actors and political parties contributes to domestic terrorism but the effect is mediated by political polarization and tribalism within countries (Piazza, 2021). According to rather similar results in a different paper, discussions on social media between opposing groups are associated with offline physical violence (Gallacher et al., 2021). These results further align with those showing that disinformation spread by foreign actors is associated with physical violence (Arayankalam and Krishna, 2021). A relation has been also found between online hate speech and offline hate crimes (Williams et al., 2020). To this end, it has also been observed that white supremacist online propaganda tends to later appear offline in the form of fliers, banners, and graffiti (Diab et al., 2023). Closer to the present study, it has been observed that exposure to online news is positively associated with participation in political demonstrations, including particularly in authoritarian countries (Kirkizh and Koltsova, 2021). While there is thus a solid foundation for the present work, thus far, the relation between online disinformation and offline political protests has not been explored. This gap in existing research is surprising already due to the studies exploring the relation of online phenomena to violence, which is a much stronger presumption than a mere political protest.
@@ 2 Framing
The disinformation and propaganda disseminated by ISIS and Russia have had one thing common; the primary (but not the only) target has been Western democracies, and the intention has been to deepen the existing societal cleavages and sabotage the internal cohesion within Western countries by exploiting weaknesses in the current engagement economy and conducting violent offline attacks, whether via terrorism or assassinations (Dundon and Houck, 2022; Hamilton, 2019; Kalniete and Pildegovic̆s, 2021; Rosenblatt et al., 2019; Sultan, 2019; Watling et al., 2024; Zhang et al., 2021). Hate is a powerful method, whether in the hands of Muslim minorities or right-wing extremists.
To reach the goals, which, as said, are a distinct trait of disinformation, time and sufficient resources are needed. In a sense, one still metaphorically needs the printing press, the airplane, or the television channel; to exploit societal weaknesses, one needs to be well-educated. The same applies when the goal of the exploitation is to provoke political protests in foreign countries. Education, staffing, and resources are thus another factor separating state-centric disinformation from other false information propagated by laypeople. Though: if the online and offline realms intervene, so do the old logic and the new logic of media; many national disinformation outlets, some of which are linked to foreign disinformation organizations, rely on national media for their criticism and distortion of narratives (Toivanen et al., 2022). For building such disinformation outlets, nevertheless, time is needed even when the result would be amateurish due to lack of education.
@@ 3.3 Independent Variables
Six independent variables are included for capturing the policy-side. A few examples are in order to contextualize this side. Although heavily criticized (Schulz, 2020) and now deprecated, the Network Enforcement Act (NetzDG) in Germany is a good example of legislative attempts to filter online hate speech. Likewise, in 2018, a comparable law was passed in France; it allows public authorities to remove fake information or even block sites that publish such content (Nagasako, 2020). In addition to these national laws, the European Union has been actively enacting different legislative measures and informal codes of conduct, including for disinformation and hate speech (Buiten et al., 2020; Pataki et al., 2021). On the side of even more drastic measures, particularly authoritarian countries in conflict zones have increasingly resorted to Internet shutdowns; among the memorable historical examples was the 2011 shutdown of the Internet in Egypt in the face of the Arab Spring uprisings (Arnaudo et al., 2013). Thus, there is a well-founded rationale to consider also the policy-side variables.
@@ 4 Results
Intercept 2.59-2.59 4.69-4.69 0.43-0.43 2.69-2.69 4.83-4.83 0.53-0.53 2.91-2.91 4.92-4.92 0.88-0.88
Gov. disinformation: domestic 0.17-0.17 0.26-0.26 0.09-0.09 0.10-0.10 0.19-0.19 $0.00$
Gov. disinformation: abroad 0.11-0.11 0.21-0.21 0.01-0.01 0.10-0.10 0.21-0.21 $0.02$
Foreign gov. disinformation 0.07-0.07 0.16-0.16 $0.02$ 0.07-0.07 0.16-0.16 $0.02$
Party disinformation: domestic 0.13-0.13 0.21-0.21 0.05-0.05 $0.02$ 0.07-0.07 $0.10$
Party disinformation: abroad 0.10-0.10 0.18-0.18 $0.00$ $0.06$ 0.05-0.05 $0.16$
@@
Gov. filtering practice $0.05$ 0.05-0.05 $0.15$
Gov. shutdown capacity 0.18-0.18 0.28-0.28 0.07-0.07
Gov. social media alternatives $0.03$ 0.10-0.10 $0.16$
Gov. social media monitoring 0.20-0.20 0.29-0.29 0.10-0.10
Regulation of online content $0.09$ 0.01-0.01 $0.19$
@@
National online media $0.01$ 0.05-0.05 $0.07$ $0.01$ 0.05-0.05 $0.07$ 0.02-0.02 0.08-0.08 $0.05$
Regime change 0.03-0.03 0.13-0.13 $0.06$ 0.01-0.01 0.11-0.11 $0.08$ $0.01$ 0.09-0.09 $0.11$
@@
Intercept 4.87-4.87 14.05-14.05 $3.25$ 4.33-4.33 12.22-12.22 $2.87$ 2.55-2.55 14.12-14.12 $7.75$
Gov. disinformation: domestic $0.13$ 0.06-0.06 $0.32$ $0.47$ $0.17$ $0.76$
Gov. disinformation: abroad 0.74-0.74 1.04-1.04 0.43-0.43 0.61-0.61 0.99-0.99 0.21-0.21
@@
Party disinformation: domestic $0.07$ 0.11-0.11 $0.25$ $0.61$ $0.34$ $0.88$
Party disinformation: abroad 0.80-0.80 1.06-1.06 0.56-0.56 0.62-0.62 0.96-0.96 0.28-0.28
Gov. filtering capacity $0.12$ 0.34-0.34 $0.59$
Gov. filtering practice 0.51-0.51 0.94-0.94 0.06-0.06
Gov. shutdown capacity 0.05-0.05 0.56-0.56 $0.48$
@@
Gov. social media monitoring 0.38-0.38 0.65-0.65 0.10-0.10
Regulation of online content 0.38-0.38 0.65-0.65 0.10-0.10
@@
Regime change 0.28-0.28 3.65-3.65 $3.10$ 1.08-1.08 3.94-3.94 $1.79$ 2.15-2.15 6.52-6.52 $2.15$
$\ln(\textmd{population})$ $0.47$ $0.00$ $1.00$ $0.59$ $0.19$ $1.05$ $0.45$ 0.16-0.16 $1.14$
$\ln(\textmd{GDP per capita})$ 0.15-0.15 0.45-0.45 $0.15$ 0.36-0.36 0.66-0.66 0.07-0.07 0.28-0.28 0.60-0.60 $0.03$
@@
Intercept 0.98-0.98 4.84-4.84 $1.79$ 18.46-18.46
Gov. disinformation: domestic $0.41$ 0.35-0.35 0.14-0.14 $0.24$
Gov. disinformation: abroad 0.47-0.47 $0.04$ $0.58$ 0.48-0.48
Foreign gov. disinformation $0.53$ 0.28-0.28 <0.01<0.01 $0.53$
Party disinformation: domestic 0.47-0.47 $0.16$ 1.25-1.25 $0.48$
Party disinformation: abroad $0.51$ 0.29-0.29 1.06-1.06 0.98-0.98
Gov. filtering capacity 0.06-0.06 $0.15$ $0.15$ 0.69-0.69
Gov. filtering practice 0.18-0.18 $0.17$ $0.68$ $0.07$
Gov. shutdown capacity 0.95-0.95 0.03-0.03 $0.32$ $0.82$
Gov. social media alternatives $0.29$ 0.04-0.04 0.65-0.65 $0.55$
Gov. social media monitoring 0.60-0.60 0.07-0.07 0.21-0.21 0.38-0.38
Regulation of online content 0.71-0.71 $0.36$ 0.25-0.25 0.45-0.45
National online media $0.16$ 0.03-0.03 0.11-0.11 0.01-0.01